A total of $12,000 is invested in two funds paying 3% and 3.4% simple interest. The annual interest is $378. How much is invested in each fund?
Can you show the steps as well so I can figure out how to do it myself? Thank you!

Answer by mananth(16946)

Fund A 3.00% per annum ---x
Fund B 3.40% per annum ---y

x + y= 12000 ------------------------1
3.00% x + 3.40% y= = $378.00
Multiply by 100
3 x + 3.4 y= = $37,800.00 --------2

Multiply (1) by -3
we get

-3 x -3 y= = -36000.00

Add this to (2)

0 x 0.4 y= = $1,800.00

divide by 0.4

y = $4,500.00 investment at 3.40%
Balance $7,500.00investment at 3.00%
CHECK
$7,500.00 --------- 3.00% ------- $225.00
$4,500.00 ------- 3.40% ------- $153.00
Total -------------- $378.00
m.ananth@hotmail.ca